OpenStudy (anonymous):

Please help me... I really need help The Miller's are comparing babysitting services for their two children. The Babysitter's Block charges a rate of $15 an hour per child plus an initial fee of $25 per household. Super Sitters charges a flat rate of $20 an hour per child. For _______ hours of service, both babysitting services would cost the same.

OpenStudy (anonymous):

So you would set this problem up like so: 15x+25 = 0 for the original Babysitter's Block, and then 20x = 0 for the Super Sitters. You would then set these equations equal to each other: 15x+25 = 20x. Subtract 15x from both sides, and you will have 5x = 25. Solve for x, and you have the amount of hours.

OpenStudy (anonymous):

im sorry im confused i've been doing homework for over 7 hours today and my brain is dead... everything is due tomorrow... is there another way to explain it? @armyengineer225

OpenStudy (anonymous):

i would probably normally be able to do this...

OpenStudy (anonymous):

Well, think of it like this: You have a set rate of $15 dollars per hour and $20 per hour for each job, right? So these wouldn't normally change unless you multiplied by the amount of hours (x). For the first block, you have an initial fee of $25 per household, so this amount is just added on to the final amount (it's not affected by how many hours is worked). Does that help?

OpenStudy (anonymous):

yes thank you :)

OpenStudy (anonymous):

No problem, hope you finish your math!

OpenStudy (anonymous):

thnk you again... do you think if i need more help you can help me? @armyengineer225

OpenStudy (anonymous):

Definitely, I'll help whenever I can.

OpenStudy (anonymous):

Which of the following equations has exactly one real solution? A. 8(x + 12) = 8x + 96 B. 8x + 12 = -8x - 17 C. 8x + 12 = 8x - 17 D. 8x + 12 = 8x + 12 can you help me with this? @armyengineer225

OpenStudy (anonymous):

is it b?